OpenStudy (anonymous):

whats the inverse function of y=x^2+9x < and equal to 0

OpenStudy (anonymous):

isn't it when you change x to y and y to x?

OpenStudy (anonymous):

oh yeah it iss

OpenStudy (anonymous):

okay then, start replacing y with x and x with y :)

OpenStudy (anonymous):

you will end up with: x= y^2 + 9y = y(y+9) so y<= 0 and y >= -9 I guess.

OpenStudy (anonymous):

yes it is ! thankss again

OpenStudy (anonymous):

np :)